Grade 8Algebra

a father is 30 years old as his son,and one year ago he was 4 times as old as his son.find their present ages.

thaniyaathee
9 Years agoGrade 8

1 Answer

Tanishq Chourishi
9 Years ago
let's take the age of father = x
and the age of son = y
then as per the above condition 
x-y=30.....(1) 
and x-1=4(y-1)
x-4y=-3.......(2)
 
(1)-(2) we get
-3y=-33  ==>y=11
then x=41
thus the present age of father and his son is 41 and 11 respectively......